# Dual formulation of the maximum entropy method applied to analytic continuation of quantum Monte Carlo data

### v1

@@ 5 Conclusions
We have compared our dual Newton algorithm with Bryan’s heavily cited algorithm. From these comparisons, we conclude that there are two advantages of our approach. First, unlike Bryan’s algorithms we search for the best positive normalization of the estimate rather than force the estimate to match the default model’s positive normalization. We argue that this capability is important for the periodic Laplace transform (i.e., finite temperature systems). Second, our approach optimizes in a reduced search space without sacrificing singular basis vectors. This addresses Rothkopf’s results which showed that removing singular vectors prevents the MEM from representing sharply peaked functions at large $\omega$ [9]. In the results section, we demonstrate that when the noise is small both algorithms yield the same mean and variance estimates, but in presence of large noise our algorithm differs from Bryan’s algorithm. Knowing that Bryan’s algorithm makes an assumption that the gradient of the fidelity term lies in the singular space while our algorithm does not, these differences suggest that noise also violates Bryan’s assumption. In summary, Bryan’s algorithm relies on assumptions that the noise is small and the desired function is sharply peaked at small $\omega$. As a result we echo Rothkopf’s statement, “either systematically extend the search space within the MEM or abandon the MEM in favor of one of the many modern Bayesian approaches developed over the past two decades.” We have expanded the search space within MEM and in doing so we have not seen the introduction of prominent ringing artifacts.
Future applications of this method are broad; the dual MEM routine used in this work can handle a generic kernel, so it can be readily applied to the applications mentioned in the introduction section. For the plasma physics community, future work could extend the cost function and solve for all of $S(k,\omega)$ simultaneously, rather than solving each $k$ individually. This same sentiment is shared by the Lattice QCD community, which wants to estimate spectral functions $\rho(\omega,T)$ for all temperatures simultaneously rather than a given $T$. In both cases, solving for all values simultaneously would enhance smoothness across the additional variable. Additionally, for the plasma physics community this work has shown that entropic regularization can invert the Laplace transform, which matches sentiments of other fields and opens the possibility of conducting instrument deconvolution using the Laplace transform instead of the Fourier transform. Lastly, future work may wish to improve over the Newton-conjugate gradient method discussed in A or explore alternatives to using the discovered positive normalization of $x$ for the default model $\mu$.
@@ Acknowledgments
Thomas Chuna thanks Matt King-Roskamp for his expertise and insightful discussions regarding what constitutes a proper problem formulation in mathematics. This work was partially supported by the Center for Advanced Systems Understanding (CASUS) which is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Union’s Just Transition Fund (JTF) within the project Röntgenlaser-Optimierung der Laserfusion (ROLF), contract number 5086999001, co-financed by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2022 research and innovation programme (Grant agreement No. 101076233, ”PREXTREME”). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The PIMC calculations were partly carried out at the Norddeutscher Verbund für Hoch- und Höchstleistungsrechnen (HLRN) under grant mvp00024, and on a Bull Cluster at the Center for Information Services and High Performance Computing (ZIH) at Technische Universität Dresden. M.P. Friedlander is partially supported by a grant from the Natural Science and Engineering Council of Canada.

### v2

@@ 5 Conclusions
We have compared our dual Newton algorithm with Bryan’s heavily cited algorithm. In section 3, we discuss two theoretical advantages of our approach. First, unlike Bryan’s algorithms we search for the best positive normalization of the estimate. We argue that this capability is important for the periodic Laplace transform (i.e., finite temperature systems). Second, our approach optimizes in a reduced search space without sacrificing singular basis vectors. This addresses Rothkopf’s results which showed that removing singular vectors prevents the MEM from representing sharply peaked functions at large $\omega$ [9]. In section 4, we demonstrate that when the noise is small both algorithms yield the same mean and variance estimates, but in presence of large noise our algorithm differs from Bryan’s algorithm. Knowing that Bryan’s algorithm makes an assumption that the gradient of the fidelity term lies in the singular space while our algorithm does not, these differences suggest that noise also violates Bryan’s assumption. In summary, Bryan’s algorithm relies on assumptions that the noise is small and the desired function is sharply peaked at small $\omega$. As a result we echo Rothkopf’s statement, “either systematically extend the search space within the MEM or abandon the MEM in favor of one of the many modern Bayesian approaches developed over the past two decades.” We have expanded the search space within MEM and in doing so we have not seen the introduction of prominent ringing artifacts.
Future applications of this method are broad; the dual MEM routine used in this work can handle a generic kernel, so it can be readily applied to the applications mentioned in the introduction section. For the plasma physics community, future work physics-oriented work could apply this method to other number densities and temperatures to observe how the roton feature varies with these parameters. Future methods-oriented work could extend the cost function and solve for all of $S(k,\omega)$ simultaneously, rather than solving each $k$ individually. The same sentiment is shared by the Lattice QCD community, which wants to estimate spectral functions $\rho(\omega,T)$ for all temperatures simultaneously rather than a given $T$. In both cases, solving for all values simultaneously would enhance smoothness across the additional variable. Additionally, for the plasma physics community this work has shown that entropic regularization can invert the Laplace transform, which matches sentiments of other fields and opens the possibility of conducting instrument deconvolution using the Laplace transform instead of the Fourier transform. Lastly, future work may wish to improve over the Newton-conjugate gradient method discussed in A or explore alternatives to using the discovered positive normalization of $x$ for the default model $\mu$.
@@ Acknowledgments
This work was partially supported by the Center for Advanced Systems Understanding (CASUS) which is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Union’s Just Transition Fund (JTF) within the project Röntgenlaser-Optimierung der Laserfusion (ROLF), contract number 5086999001, co-financed by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2022 research and innovation programme (Grant agreement No. 101076233, ”PREXTREME”). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The PIMC calculations were partly carried out at the Norddeutscher Verbund für Hoch- und Höchstleistungsrechnen (HLRN) under grant mvp00024, and on a Bull Cluster at the Center for Information Services and High Performance Computing (ZIH) at Technische Universität Dresden. M.P. Friedlander is partially supported by a grant from the Natural Science and Engineering Council of Canada.